Shift Summary for 9/3/99 Evening Shift

Vivian O'Dell ((no email))
Fri, 3 Sep 1999 23:57:30 -0500

Shift summary for EVE shift, 9/03/99

On Shift: A. Alavi, V. O'Dell

Hours of beam available at all intensities: 3.5h
Hours of beam available at nominal intensity: 3.5h
Hours of E832 data (special runs not included): 3.5h
Hours of special runs (laser scans, muon runs...):0h

Accounting of lost time: Quench, fridge problems, low beta, all those
words I don't understand.
========================================================================
Runs ended this shift:

14425 - 12.2 K vac 2pi0 stopped when tapes nearly full, no beam.

Tapes brought to FCC this shift: QK(S,C,K)375-377, QKB132
Accesses during shift:
========================================================================
Overview of shift: Is Labor Day weekend a good time to collect data? Is
this what it will be like during the Christmas
"shutdown" ?

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14425 15:30 Evening shift turns up.

14425 15:45 Day shift leaves after putting Blondie on the CD
player, promising to return with beer and disco
lights.

14425 15:49 BobT turns up. Evening shift nervously hides disco
lights & beer.

14425 53 15:50 Beam back after a long hiatus.

14425 57 15:55 Valeri (who will be on owl shift) notices that
beam is aborting every other spill.

14425 62 16:00 Viv finally agrees that beam is aborting every
other spill.

14425 16:05 Jason Hamm, hearing about the party, arrives. Spotting
BobT, he nervously pretends to work.

14425 ~68ish 16:10 No beam. F17 kicker problems. 10 minute estimate.

14425 16:18 Hogan, hearing about the party, arrives. He also
nervously pretends to work.

14425 BobT starts wandering around. We all get nervous.
He doesn't spot the beer or the disco lights, but
he does notice stacks and stacks of labelled tapes,
courtesy of the day shift. Hogan starts talking about
volleyball.

14425 104 16:55 Beam back. Message on screen claims beam will
be intermittant due to continuing F17 kicker
problems, but no problems for the first 2 spills.

14425 120 17:25 Everyone seems to have left. No party tonight. One
shift crew member wanders off to take tapes to
FCC.

14425 18:25 Shift crew member finally returns, subway sandwich
in hand and stinking of Starbucks coffee. Where did
those tapes go anyway?

14425 19:00 Heated argument between one former KTeV member (now
a Columbia postdoc) and a current KTeV member (now
a FNAL Wilson Fellow) over the definition and meaning
of radiation length. Definately no party tonight.
Shift crew gets very depressed and puts Tracy Chapman on the CD player.

14425 19:10 The two arguing physicists leave. Shift crew sinks
into depression.

14425 ~218 19:30 Quenchola. Estimate, 1 hour. Since tapes are nearly
full (> 90%), we figure this is a good time to
swap them. In order not to waste one precious
second of beam time, we spend 30 minutes to
develop a well-thought out tape swapping plan,
complete with resource loaded WBS numbers.

20:05 While one shift member was worrying about the
lack of resources at level 5, the run was stopped
and the tapes were miraculously changed. If only the
collider experiments knew our secrets, they'd be
ready to take data now.

14426 20:10 So tired, tired of waiting, tired of waiting for
beam...

20:45 Still waiting for beam. Channel 13 updated, low beta
problems, new estimate: no estimate.

14426 ~100 22:35 Beam to users!

14426 115 23:00 After a couple of aborts, beam gone again...
Shift crew sinks back into depression.
"C4 fridge problems; investigating."

14425 23:30 Valeri shows up for the owl shift. He demands
to do TRD HV scan and the depressed shift
crew is powerless to stop him.

========================================================================
Follow up Notes on Previous Shift:
o kept our eyes on channel 226.

========================================================================
Items to Follow up on Next Shift:
o keep your eyes on channel 226.
o also 961 has (infrequent) cap id problems.

========================================================================